### v2

Abstract
[Abstract]This paper presents the development of a Physics-realistic and Photo-realistic humanoid robot testbed, PR2, to facilitate collaborative research between Embodied Artificial Intelligence (Embodied AI) and robotics. PR2 offers high-quality scene rendering and robot dynamic simulation, enabling (i) the creation of diverse scenes using various digital assets, (ii) the integration of advanced perception or foundation models, and (iii) the implementation of planning and control algorithms for dynamic humanoid robot behaviors based on environmental feedback. The beta version of PR2 has been deployed for the simulation track of a nationwide full-size humanoid robot competition for college students, attracting 137 teams and over 400 participants within four months. This competition covered traditional tasks in bipedal walking, as well as novel challenges in loco-manipulation and language-instruction-based object search, marking a first for public college robotics competitions. A retrospective analysis of the competition suggests that future events should emphasize the integration of locomotion with manipulation and perception. By making the PR2 testbed publicly available at https://github.com/pr2-humanoid/PR2-Platform, we aim to further advance education and training in humanoid robotics.
Video demonstration: https://pr2-humanoid.github.io/
keywords:
embodied AI, humanoid robots, simulation, education robots1 Introduction
Recent academic and industrial advancements in humanoid robotics have revealed critical gaps in educational infrastructure and research accessibility. While miniature humanoid robots like DARwIn-OP ha2011development and NAO gouaillier2008nao have served as entry-level educational tools, their servo-based actuation system and limited payload capacities restrict investigations into advanced topics such as whole-body control, dynamic locomotion, and manipulation. Although emerging force/torque controlled platforms attempt to address these limitations grimminger2020open, liu2022design, their high maintenance requirements and substantial costs still impede widespread adoption.
Simulation environments, while not perfectly replicating the real world, have emerged as cost-effective alternatives for robotics education and research, demonstrating effectiveness across various domains carlone2022visual, moon2024autonomous. However, there is a significant gap in the availability of simulation platforms specifically designed for diverse topics in full-size humanoid robots. Mainstream robotics platforms including Gazebo koenig2004design, Webots michel2004cyberbotics, V-REP rohmer2013v, MuJoCo todorov2012mujoco, and PyBullet coumans2021pybullet, prioritize high-fidelity dynamics simulation at the expense of visual realism, thereby hindering integrated perception-action studies. Conversely, simulation environments derived from gaming engines, such as VRGym xie2019vrgym, ManiSkill mu2021maniskill, OmniGibson li2023behavior, Habitat puig2023habitat, and RoboCasa nasiriany2024robocasa, achieve photorealistic rendering but oversimplify physical interactions through kinematic approximations. For instance, they often assume that the robot can precisely achieve the desired pose and that the environment does not interact with the robot in unexpected ways, overlooking the fundamental challenge in humanoid robots–dynamic whole-body motions. This fundamental dichotomy creates a critical void in platforms capable of simultaneously supporting dynamic locomotion studies and vision-based manipulation tasks, as illustrated in Table 1.
The development landscape further compounds these challenges, as evidenced by high-profile competitions like the DARPA Robotics Challenge (DRC) johnson2017team and RoboCup. While these initiatives drive technological innovation, their reliance on custom-built hardware solutions and requirement for specialized expertise effectively exclude undergraduate participation and limit knowledge transfer. As a result, participants in these competitions have primarily been postgraduate students, professional engineers, and researchers.
To bridge these complementary gaps in simulation capabilities and educational accessibility, we present PR2, a novel humanoid robot testbed integrating three critical innovations: (1) Co-simulation architecture combining rigid-body dynamics with real-time photo-realistic scene rendering; (2) Modular control framework supporting hierarchical whole-body control and sensorimotor integration; (3) Multi-modal interface for foundation model integration. The efficacy of the platform was validated through a national collegiate competition involving 137 teams (402 participants) executing six progressively challenging tasks that span bipedal locomotion, multimodal manipulation, and interaction with vision and language foundation models, as shown in Figure 1. The four-month competition provided insight into the challenges students faced, deepened our understanding of humanoid robot education, and inspired future improvements in the affordability and accessibility of humanoid robot research and education. Following a detailed report on the main results and findings from the competition, we enhanced our testbed and made it publicly available.
| Feature |
PR2 |
Gazebo |
MuJoCo |
Webots |
V-REP |
PyBullet |
ManiSkill |
Habitat |
OmniGibson |
RoboCasa |
VR-Gym |
| Physics Accuracy | ++ | + | ++ | + | + | + | − | − | + | + | − |
| Photorealism | ++ | − | − | − | −− | −− | ++ | ++ | + | ++ | + |
| Environment Creation | ++ | −− | −− | − | −− | − | + | ++ | ++ | ++ | + |
| Model Training | + | − | + | − | − | + | + | + | + | ++ | + |
| ROS Support | + | ++ | + | ++ | ++ | + | −− | − | − | − | − |
| Dynamic Locomotion | ++ | ++ | ++ | + | + | + | −− | − | − | − | −− |
| Loco-Manipulation | + | + | ++ | − | − | − | + | − | + | + | − |
| Vision-Language Integration | ++ | −− | − | − | −− | − | ++ | + | + | + | + |
| Real-Time Co-Simulation | + | ++ | ++ | ++ | ++ | ++ | + | − | + | − | ++ |
| Humanoid Controller Integration | + | − | − | − | − | − | − | − | − | − | − |
| Competition Scene & Tasks | ++ | + | −− | −− | −− | −− | −− | − | − | − | −− |
Key contributions of this work include:
Open-source release of a unified simulation environment for humanoid robot education and research, addressing the realism-dynamics trade-off.
Curriculum framework for learning in humanoid robot perception, planning, control, and AI model integration.
Quantitative analysis of educational barriers in humanoid robotics through competition metrics.
1.1 Related Work
1.1.1 Virtual Reality
Contemporary virtual reality platforms have become foundational tools for developing robotics and AI simulators, leveraging advanced rendering engines to achieve photorealistic environments. Unreal Engine-based solutions such as CARLA dosovitskiy2017carla for autonomous cars, AirSim shah2018airsim for aerial vehicles, VRGym xie2019vrgym, and VRKitchen gao2019vrkitchen for robot learning demonstrate this trend, while Unity-powered platforms like AI2-THOR kolve2017ai2, VirtualHome puig2018virtualhome, and ThreeDWorld gan2020threedworld showcase comparable capabilities. These systems prioritize visual fidelity over physical verisimilitude, employing kinematic approximations for robot-object interactions. This architectural limitation becomes particularly problematic when simulating compliant contacts or inertial effects – critical factors in humanoid locomotion attempts to integrate basic physics engines. These findings underscore a fundamental divergence in design priorities: gaming-derived environments optimize for visual immersion, whereas robotic systems require precise dynamics resolution.
1.1.2 Dynamic Interaction in Simulation
The robotics community has witnessed growing adoption of NVIDIA Isaac Sim as a versatile platform for locomotion and manipulation studies, despite its utilization of conventional physics engines (e.g. Bullet or PhysX) shared with VR systems. This trend is exemplified by OmniGibson li2023behavior – the latest evolution of the Gibson environment xia2018gibson, xia2020interactive, and emerging Embodied Artificial Intelligence (Embodied AI) benchmarks including RoboCasa nasiriany2024robocasa, ARNOLD gong2023arnold, and ORBIT mittal2023orbit. While these platforms facilitate basic physics-based interaction modeling, they typically focus on kinematics rather than dynamics—a critical aspect for full-size humanoids. The proposed PR2 testbed enables full dynamic simulation for aggressive motions and manipulation with external objects, a key direction for future research and applications in humanoid robotics. Additionally, we have incorporated essential tools and algorithms of humanoid robots for beginners to enhance the platform’s accessibility.
1.2 Overview
We organize the remainder of the paper as follows. In Section 2, we introduce the system architecture of the PR2 testbed, detailing its key modules. Section 3 presents the competition setup based on the PR2 testbed and describes the customizations made to support a wide range of humanoid robot tasks. Section 4 highlights the capabilities of PR2 and summarizes the competition results. Finally, in Section 5, we conclude the paper with a discussion on the future directions for the humanoid testbed.
2 System Architecture
The system architecture of PR2 is illustrated in Figure 2. PR2 is built on top of NVIDIA Isaac Sim, which leverages advanced GPU-enabled graphics and physics simulation with Nvidia PhysX 5, enabling users to investigate and test robotic skills in a physics-realistic and photo-realistic simulation environment. Designed to be modular, PR2 simplifies workflows in interacting with humanoid robots with an easy-to-use interface for users who can incorporate their own modules into the platform. By extending the core components of NVIDIA Isaac Sim, PR2 consists of three main modules described in the following subsections.
2.1 Task Module
Leveraging abundant digital assets in the form of open-source Universal Scene Description (USD) files, PR2 offers a fully interactive task environment with diverse indoor and outdoor spaces populated with various objects, moving beyond the limitations of static meshes. To achieve high rendering quality, realistic ray-traced ambient light, and other lighting effects are integrated into the scene, significantly enhancing the fidelity of simulated visual sensor data. As a result, the platform can support tasks such as visual navigation, object detection, and other perception-related activities by capturing high-fidelity RGB and depth information from the robot’s egocentric view. For object interactions, PR2 employs a contact sensor mechanism to determine the contact configurations between the robot and the target object.
2.2 Physics-based Simulation
To interface with the virtual humanoid, we adopt the API conventions of Nvidia Isaac Sim. The Robot class is designed to load the robot model from a USD file and manage the associated physics handles required for setting and reading simulation states. For robot joint actuation, we offer three control modes: position, velocity, and torque. In the torque-controlled mode, users can directly apply torque commands to the robot, which are interpreted as joint efforts. In the position- and velocity-controlled modes, users specify commands for joint positions or velocities, which are then converted into joint efforts using internally implemented spring-damper models. Beyond robot actuation, we incorporate rigid body collision handling through PhysX. PR2 can also support sophisticated physical simulations, including object deformation and fluid dynamics gong2023arnold. This setup not only facilitates accurate robot control but also enhances the realism of interactions within the simulated environment.
2.3 Robot Controller
Due to the underlying physics-based simulation, the humanoid imported into PR2 would not be dynamically stable by itself. However, developing dynamic controllers of bipedal humanoids from scratch could be prohibitively difficult for beginners in this field. As a result, providing essential built-in controllers could expedite users’ knowledge of such fields before implementing their own or investigating other topics on top of it.
We developed basic walking and jumping controllers in Drake drake, which receive user action commands, generate feedback signals from the simulator, and return the effort value of each leg joint that can be executed by the robot to the simulator through TCP/IP communication. The default parameters for these controllers would support basic stabilization and tracking but are not ideal. Users are encouraged to modify the controllers to obtain more stable and more responsive execution, and they have the option to replace these controllers with custom solutions for more sophisticated and specialized control.
An example action command is shown in Figure 3(a) where and are the number of arm and leg joints. The text in the comment indicates the type of data expected and the possible values that can be assigned to a specific field. Specifically, users can choose to actuate each robot joint by position, velocity, or effort (i.e.torque) based on the provided joint values array. Users can also empirically adjust the corresponding stiffness and damping for each joint, which influences the actual joint efforts applied in position- or velocity-controlled modes. For manipulation tasks, symbolic pick and release actions are designed. The pick action, triggered by specifying either the left or right hand, attaches the target object to the robot’s end-effector if the distance between them is within a certain range. The object detaches from the robot when it receives a True signal for release.
Once an action is executed by the robot, an observation would be generated and returned to the user. An example of the observation space is provided in Figure 3(b), where and are the height and width of camera images. The observation data structure provides a comprehensive representation of the robot’s state, including its joint positions, velocities, and applied efforts for both arms and legs. Additionally, the robot’s body poses and its linear and angular velocities w.r.t. the world frame are available under the body_state key. The robot’s egocentric perceptual data is accessible through the cam key, which includes both the RGB image and the distance to the image plane (i.e.depth). The language_instruction key enhances support for vision-language-action tasks. A boolean value under the pick key indicates whether an object has been successfully attached to the robot’s end-effector, a critical indicator for manipulation tasks. Additional parameters, such as stiffness and damping, or any ad-hoc specifications, can be introduced by adding new keys to the observation structure, allowing for flexible and task-specific customization.
Finally, the users are responsible for developing their high-level planners that strategize the robot’s path and actions to accomplish given tasks with the required efficiency and effectiveness.
3 Competition Overview
After developing the necessary modules, we organized a nationwide collegiate humanoid robot competition as a pilot study to validate the PR2 testbed through an online virtual challenge. The competition ultimately attracted 137 teams, each with up to four students. This section details the competition setup and evaluation rubric.
3.1 Competition Trials
To comprehensively address humanoid locomotion, manipulation, and AI integration while maintaining pedagogical accessibility, we designed six progressively challenging tasks; see Table 2 for a detailed description. Figure 4 visually summarizes these tasks.
Task 1: As shown in the upper left of Figure 4, the robot is initialized randomly within the red rectangular region with a heading angle deviation within . The objective requires the robot to traverse uneven terrain and reach the green circular target zone on the left. The task is considered successful if the robot’s body moves into the circle region within . The robot body and joint states, as well as start position and orientation, are provided as feedback, with the goal center position provided as prior information. This task tests the robustness of walking control as a basic of the humanoid robot.
| Task Specification | Scoring Rubric | Feedback | Require Module | |
| Task 1 (15 points) Uneven terrain walking | goal center position | walk into the task region (+5) | robot body & joint states start position & orientation | walking velocity planner |
| walk half long (+5) | ||||
| walk into finish region (+5) | ||||
| Task 2 (15 points) Valve turning | valve center position valve dimensions valve ID | walk into the task region (+5) | robot body & joint states start position & orientation arm pick state | loco-manipulation planner and controller |
| turn the valve (+5) | ||||
| turn the valve more than (+5) | ||||
| Task 3 (15 points) Vertical jumping | moving bar ID | step over the moving bar (+10) | robot body & joint states start position & orientation moving bar position & velocity | jumping motion planner |
| jump over the moving bar (+15) | ||||
| contact with the moving bar (-5) | ||||
| Task 4 (10 points) Slope climbing | goal center position slope incline angle | walk into the task region (+5) | robot body & joint states start position & orientation | slope walking controller |
| walk into the finish region (+5) | ||||
| Task 5 (20 points) Step climbing and button pressing | button center position step dimensions | walk onto the step (+5) | robot body & joint states start position & orientation | step walking planner walking velocity planner arm motion planner |
| walk over all the steps (+5) | ||||
| touch the button (+5) | ||||
| touch the button with arm (+5) | ||||
| Task 6 (25 points) Indoor object search and transport | 3D model of 9 objects coffee table center position a sentence of task: put the object on the coffee table | walk to the pick region (+5) | robot body & joint states start position & orientation arm pick state camera RGB-D information | loco-manipulation controller Vision-Language-Action model |
| pick up target object (+10) | ||||
| walk to the table (+5) | ||||
| place the object on the table (+5) |
Task 2: Illustrated in the middle left of Figure 4, the robot needs to turn the red valve to more than within to be considered successful, which is a loco-manipulation task. For simplification of contact dynamics (details provided in Section 3.3), a fixed joint is automatically established between the robot’s end-effector and the valve when their proximity falls below a threshold, by sending the pick command.
Task 3: The center plot of Figure 4 illustrates the task specifications, where a bar moves horizontally from right to left, requiring a jumping movement of the robot over the bar and landing stably within to acquire full points. The position and velocity of the bar are provided as feedback for jumping motion planning and control. This task tests the capability of humanoid robots to conduct highly dynamic motions.
Task 4: The lower left of Figure 4 demonstrates the initialization area and task goal region of this task. The robot needs to walk through a slope within , with the provided slope incline angle as prior.
Task 5: The robot needs to first walk through both steps and the plain floor and then contact the red button within to accomplish this task (see lower right plots of Figure 4). The dimensions and number of steps are provided as the task prior.
Task 6: This is a comprehensive indoor object search task (lower right of Figure 4), where a sentence of the task will be provided at first to describe the task. Three possible objects are designed: the cup on the kitchen table shown on the upper left, the bear on the table in the playroom (mid bottom), and the cube on the desk of the studyroom (right bottom), these objects are randomly initialized in the rectangular region of each room. The robot needs to find the assigned object and translate it to the table in the living room within . The 3d model of each object is provided as prior, and the camera view of the robot is streamed out as feedback.
These tasks are operated independently rather than sequentially, which ensure that participants who may struggle with one aspect still have opportunities to engage with other challenges. For example, difficulties in manipulation in Task 2 do not impact the performance in subsequent tasks involving locomotion, such as jumping (Task 3) and climbing slopes and stairs (Task 4 and 5). For each task, if the robot falls, walks outside the task region, or reaches the assigned time limit, the simulation will terminate automatically.
| Dimension Parameters | ||||
| Total mass | Total height | Pelvis width | Thigh length | Calf length |
| 34.5 [kg] | 1.20 [m] | 0.22 [m] | 0.23 [m] | 0.26 [m] |
| Foot length | Upper arm length | Forearm length | ||
| 0.15 [m] | 0.24 [m] | 0.20 [m] | ||
| Motion Range & Joint Peak Torque | ||||
| Hip Yaw | Hip Roll | Hip Pitch | Knee Pitch | Ankle Pitch |
| 48 [Nm] | 110 [Nm] | 110 [Nm] | 110 [Nm] | 48 [Nm] |
3.2 The Robot
The KUAVO v1.0 platform, a humanoid robot with a height of and a weight of he2024cdm, is used in the virtual competition. It incorporates 18 Degree of Freedoms, 5 for each leg and 4 for each arm. In this competition, all joint actuators of the KUAVO robot operate in the torque-controlled mode, better replicating the control architecture of a full-size humanoid. For the details of the hardware platform’s specifications, please refer to Table 3.
3.3 The Virtual Testbed
The PR2 utilizes the Articulations package from the Isaac Sim Core extension to achieve realistic dynamic behaviors through whole-body torque control on the KUAVO robot. The Articulations package provides a high-level interface for managing robots using the Root Articulation API, facilitating effective handling of their attributes and properties.
To ensure stability during simple standing as well as complex tasks such as traversing rough terrains or performing loco-manipulation, three basic controllers are provided for the KUAVO during the competition: (i) a standard walking controller that accepts body velocity commands (3-DoFs) for Tasks 1, 2, 4, and 6; (ii) a dynamic jumping controller based on a centroidal dynamics model he2024cdm, which takes inputs for jump time, velocity, and height for Task 3; and (iii) a step walking controller based on the Zero Moment Point (ZMP) method scianca2020mpc for Task 5. The arm controller is not provided and must be developed by the teams to complete Tasks 2, 5, and 6 successfully. Critical implementation details for each task are outlined below:
Task 1: The surface of the -meter-long rocky road features fine-grained collision geometry that results in unstable footsteps with different contact forces applied to the robot.
Task 2: The users could turn the valve (diameter around ) either by hitting it with the robot arm’s distal end or by establishing a fixed attachment with the valve when the distal end reaches the enlarged valve volume (see Figure 5(b)) by sending a “pick” command; this allows the valve revolves together with the robot’s locomotion.
Task 3: Users can either design a gait that steps over the bar and adjust the built-in walking controller to track the gait stably or use the provided jumping motion planner to jump over it.
Task 4: This task is similar to Task 1, but the users need to tune the planner and controller to climb the slope with a incline.
Task 5: Users must design a gait for the robot to ascend stairs ( high and wide) while maintaining the desired heading direction to press a button. The pressing mechanism is implemented using the PhysX Contact Report API to detect if the robot (preferably its arm) makes contact with the button’s surface.
Task 6: The robot would need visual perception (to recognize scenes and objects), language understanding (to translate questions and instructions into actions), and navigation in complex environments (to move and find things). When the user sends a pick command, a fixed joint will be created only if the Euclidean distance between the end-effector and the object to be picked is less than , expressed as a spherical region as shown in Figure 5(a).
For each task, the robot’s initial pose is randomly sampled within a designated start region. Although the objective interactions in task 2 and task 6 have been simplified for the college competition, the underlying PhysX engine is fully capable of supporting more advanced manipulation tasks. For example, complex interactions, such as grasping objects with a multi-finger gripper, can be achieved by utilizing finer-grained collision-checking and more sophisticated control strategies. It is important to note that while these capabilities exist within our simulation framework, the explicit demonstration remains one of our ongoing efforts in future PR2 releases.
4 Results
The beta version of PR2 was released to participants in January 2024, with the program submission deadline set for May 20, 2024. The competition ultimately received 51 valid submissions. In this section, we first highlight three unique features of PR2. Then, we report and analyze the competition results. Finally, we summarize our key takeaway from the competition with critical lessons learned.
4.1 Platform Capability
4.1.1 Kinodynamic Locomotion
Leveraging the underlying PhysX engine and the Isaac Sim APIs, PR2 enables the execution of generated kinodynamic motion plans. With built-in controllers, Figure 6(a) shows the robot performing vertical jumping to a moderate height and to an extreme height that is pushing the controller to its limit. Such jumping trajectories are naturally kinodynamic due to the launching, flight, and landing phases during the jump all pose unique challenges in dynamics planning and control, and all require high-fidelity dynamical simulation to capture and reflect such processes.
Specifically, (i) after generating trajectories for the center of mass (CoM) and centroidal angular momentum (CAM) within the hardware capabilities, PR2 needs to accurately compute the robot’s dynamical state evolvement. (ii) In the flight phase, the platform must determine the controller’s trajectory tracking performance to reflect the disturbances from execution errors. (iii) As the robot lands from a higher jump, it experiences larger impact forces. Thus, PR2 must not only reflect different landing impacts but also propagate such impacts to the landing controller to determine robot stability. Figure 6(a) demonstrates the execution of two significantly different jump trajectories, wherein the red one takes longer for the controller to stabilize the robot after landing. The capability to reflect the subtle differences in locomotion is a critical aspect of a humanoid testbed.
4.1.2 Physics-based Interaction
Together with the realistic simulation of robot locomotion, PR2 further enables humanoid’s physics-based interaction with the environment during manipulation. Similar to the jumping example, Figure 6(b) shows how the reactive force caused by the target valve influences robot motions in loco-manipulation. With a naive controller, the reactive force causes a dramatic torso pitch, whereas a more stable controller could successfully reject the reactive force and maintain body balance.
4.1.3 VLA Interface
We introduce customized APIs that enable users to interact with large language models (LLMs) and vision-language models (VLMs), providing a new approach to exploring AI and robotics. In Task 6, users receive visual observations and language instructions to identify goals, formulate plans, and compute robot actions (e.g. joint positions or efforts) at each timestep. The target object in the language instruction is randomly selected from a set in PR2, and the robot is initialized at a random location in the same room as the target object.
Figure 7 illustrates a typical example where the instruction is put the [pencil on the desk] on the coffee table in the living room. The task involves both object detection (identifying the pencil) and scene understanding (locating the coffee table). Typically, the pencil is in a different room from the coffee table, requiring the user to generate high-level navigation and motion plans to guide the robot in avoiding obstacles and executing the necessary actions. The agent must take the pick action when it believes it is close enough to the target object. If the distance is within a specified threshold, a fixed joint is created to attach the target object to the robot’s end effector. Upon reaching the coffee table, the agent must take the release action to complete the task.
4.2 Team Performance Analysis
Figure 8 presents the score distribution percentages across the six tasks. Each bar representing a task is divided into colored segments corresponding to different score ranges.
For Task 1, the majority of participants (52.94%) achieved the full score of 15, followed by 23.53% scoring 5, and a small portion (3.92%) received a score of 10. Task 2 exhibited a more varied distribution due to the increased difficulty of balancing bipedal walking with the constraints of the contacted valve. Approximately 58.82% of participants scored between 0 and 5 (indicating an inability to turn the valve), while only 19.61% fully succeeded in this task. In Task 3, despite the provided jumping controller, nearly 70% of participants failed to score any points. As the only task involving a dynamically changing environment (i.e.the moving bar), participants struggled to design a high-level plan for the robot to jump and land at the right moments without hitting the bar. Task 4 and 5, both of which emphasized bipedal walking capabilities for climbing slopes and stairs, respectively, showed similar success rates to Task 1: 50.98% of participants in Task 4 and 37.25% in Task 5 achieved full marks. Task 6, which integrated perception, language parsing, manipulation, and walking, presented the greatest variation in scores, with only 11.76% of participants successfully completing it. The performance diversity across tasks highlighted the importance of introducing different challenges in competition.
Figure 9 further presents the individual time consumption for each score across all tasks, with each score represented by a distinct box plot that shows the distribution of times achieved by participants. The overall plot reveals varying levels of difficulty and completion times for each task. Generally, the robot requires more time as it becomes more involved in the task to achieve higher scores. Task 6 exhibits the widest range of scores and the most balanced distribution, indicating diverse levels of student performance. This task saw some students excelling, while others struggled, as evidenced by the highest proportion of participants scoring 0 and the second highest scoring 25. The presence of outliers across all tasks suggests that while most students completed within a typical range, some took significantly longer.
By categorizing the detailed score rubric into three categories—locomotion (60%), manipulation (30%), and perception (10%)—as shown in Table 2, we analyzed the average points participants scored (see Figure 10). The results indicate that participants performed relatively well in locomotion, earning nearly half of the available points (27.55 out of 60), reflecting the emphasis this area has received in humanoid literature. However, they encountered significantly greater challenges in tasks involving manipulation and perception, scoring less than 30% and 20% of the available points in these categories, respectively. Figure 11 shows some typical failure cases in manipulating the valve (Task 2) and perception (Task 6).
4.3 Lesson Learned
Throughout the competition, we hosted an online forum to address participant questions and technical issues. Combined with the competition results, this process provided valuable insights that guided our engineering efforts and may also benefit the broader robotics community in advancing research and education on humanoid robots.
4.3.1 Trilemma of rendering quality, simulation precision, and computing resource
Despite the rapid advancements in AI and robotics, access to dedicated Nvidia GPUs remains limited, particularly for students from traditional engineering backgrounds. During the competition, PR2 encountered over 20 issues related to platform launch failures, primarily due to missing or inadequate GPUs. Unfortunately, some participants were forced to withdraw due to the lack of GPU access.
In the competition, a limited number of participants had access to high-end consumer-grade GPUs, such as the RTX 40 Series, which can cost up to . To address these hardware constraints, we introduced new APIs that enable participants to adjust rendering quality and modify the simulation time-step. Reducing rendering quality can increase the frame rate, ensuring that PR2 can photo-realistically reflect the simulated physics. We conducted benchmarks on a computer equipped with a 13th Gen Intel® Core™ i9-13900KF CPU and an Nvidia 4090 GPU. With a fixed physical simulation time-step of 0.001, Table 4 indicated that the frame rate remains stable until the rendering workload approaches the system’s computational limits. Beyond a resolution of pixels, the frame rate decreases as rendering demands exceed available resources, necessitating adjustments to maintain consistent physics time-steps. Alternatively, increasing the simulation time-step can alleviate computational burdens. Larger time-steps give the physics engine coarser time slices to compute interactions, including the integration of forces and motion, leading to numerical errors and degraded fidelity. Balancing rendering quality, simulation precision, and computing resources is a crucial design consideration for promoting more accessible and equitable education opportunities in humanoid robotics.
| Resolution (pixels) | 640480 | 1024768 | 1280720 | 16001200 |
| Average FPS | 17.363 | 17.683 | 17.361 | 16.263 |
| Resolution (pixels) | 19201080 | 25601440 | 34401440 | 38402160 |
| Average FPS | 14.918 | 15.033 | 13.973 | 13.697 |
4.3.2 Model-based controller v.s. Model-free policy
Although the competition provided basic controllers for robot stabilization and tracking, some participants opted to develop and train RL-based policies instead. While both approaches can effectively generate robot motions and are encouraged, they introduce significant challenges in (i) creating a robust testbed that supports both model-based and model-free methods, potentially combining elements of them in hybrid approaches, and (ii) ensuring fair evaluation as RL often require significant computational resources for training.
4.3.3 The necessity and challenges in integrating AI-related topics
Participants generally performed well in locomotion-related tasks with the provided controllers, as shown in Figure 10. However, significant performance drops were observed when tasks incorporated elements of manipulation and visual or language perception, as reflected by lower average scores. Specifically, difficulties were most pronounced in Task 3 and 6, which required high-level planning and the integration of AI techniques. These results underscore the need for future research and education in humanoid robotics to encompass a broader range of skills and topics, addressing the challenges that arise when integrating complex AI components with robot tasks.
4.3.4 Competition with Physical Robots
The top 10 teams from the simulation competition were invited to participate in the physical humanoid robot competition held in August 2024. Five trials were designed based on the simulated tasks in PR2, with modifications to accommodate physical robots (see Figure 12). For the valve-turning and button-pressing tasks, the robot needed to locate the target object, approach it, and use its arm to either rotate a horizontal valve 180 degrees or press a button. In the uneven terrain walking task, the robot had to maintain stability while navigating rough surfaces. The door-opening task required the robot to grasp and rotate a doorknob. In the indoor object search task, an audio cue indicated the target object the robot needed to fetch, incorporating challenges in speech recognition, object detection, locomotion, and grasping. Each team had one hour for competition and participants could complete the tasks in any order, but bonus points were awarded for completing tasks consecutively.
Several sim-to-real gaps were identified during the competition. Accurately replicating the damping and friction characteristics of both the robot’s actuators and the valves used in the physical competition within the simulation environment proved challenging. This discrepancy can lead to differences in robot performance between simulated and real-world scenarios. Additionally, the simulation environment often fails to capture the complex lighting variations present in real-world settings, affecting the robot’s vision system and its ability to accurately perceive and interact with objects. Despite the substantial domain gap between simulation and real-world execution, PR2 remains an effective testbed for participants to become familiar with the robotic platform and build a foundation for the competition. On average, participants scored , , , and in the valve-turning, button-pressing, uneven terrain walking, and door-opening tasks, respectively—significantly higher than the average performance in simulation competition. However, the indoor object search task saw an average score of only , highlighting the challenges of integrating locomotion with manipulation and visual/audio perception.
5 Conclusion and Looking Forward
In this paper, we presented the development of PR2, a testbed that combines physics-realistic simulation with photo-realistic scene rendering. By integrating high-quality scene rendering with high-fidelity simulation of robot dynamics, PR2 offers a new platform for exploring full-size humanoid robot locomotion, manipulation, and perception, addressing critical aspects for future humanoid applications. The deployment of PR2 in an online collegiate competition attracted hundreds of participants, revealing the challenges junior students faced in combining manipulation and perception in humanoid robot locomotion and emphasizing the need for future research and education to focus on these complex integrations. We have released an upgraded version of the PR2 testbed, which we believe will significantly benefit future research and education in humanoid robotics, particularly for those who previously lacked access. Moving forward, our priorities include incorporating a built-in RL training scheme alongside planning and control methods, enhancing scene diversity through scene synthesis techniques yang2024physcene, and refining fine-grained grasping capabilities for loco-manipulation tasks in future PR2 releases.
Acknowledgement:
This work was supported in part by the National Natural Science Foundation of China (No. 62403064, 62403063). We thank the engineering team from Leju Robot for technical support, Ms. Zhen Chen from BIGAI for refining the figures, and all colleagues from the BIGAI TongVerse project for fruitful discussions and help with simulation developments.